OpenStudy (anonymous):

log7^n=2/3log7^8 Please solve step by step, thanks

OpenStudy (campbell_st):

ok... the left hand side can be rewritten using the index law for powers \[\log(x^a) = a \times \log(x)\] so you get \[n \times \log(7) = \frac{2}{3} \times 8\times \log(7)\] divide both sides of the equation log log(7) and you'll have your answer

OpenStudy (anonymous):

n=4 ??

OpenStudy (campbell_st):

not quite,,, won't dividing by 7 leave \[n = \frac{2}{3} \times 8\] simplify the fraction

OpenStudy (campbell_st):

oops should read dividing by log(7)

OpenStudy (anonymous):

oh my friend told me it was 16/3 but i thought that was wrong.

OpenStudy (campbell_st):

nope its the correct answer

OpenStudy (anonymous):

Oh well i solved it a different way, I thought it would be log7^8^2/3=log7^4 which equals log7^n=log7^4 which makes n = 4 ? is that method wrong ?

OpenStudy (campbell_st):

well what you really have is \[\frac{2}{3} \log 7^8 = \log(7^8)^{\frac{2}{3}}\] using the power of a power index law you get \[\log 7 ^{8 \times \frac{2}{3}}\] so n = 16/3 hope that makes sense...

OpenStudy (anonymous):

Ohhh so i had to multiply it and NOT raise by that power ?

OpenStudy (campbell_st):

well it uses a combination of index laws and log laws... so there are a few methods to solve the equation.

OpenStudy (anonymous):

Ohh ok, thanks so much for your help, I just have 1 more to do can i do it on my own and come back to you and tell me if the answer is right ?

OpenStudy (campbell_st):

just post it if you get stuck... and there are lots of people here who can help

OpenStudy (anonymous):

(2log6^4)-(1/4log6^16)=log6^x x=2 can u just tell me if this is right its my last poblem.

OpenStudy (anonymous):

btw the problem didn;t include parenthesis i thought it would just help read the problem better.

OpenStudy (anonymous):

2log6^4-1/4log6^16=log6^x is the problem

OpenStudy (campbell_st):

ok.. so the log law for subtraction is \[\log(a) - \log(b) = \log(\frac{a}{b})\] and you can rewrite the the powers \[\log(\frac{(6^4)^2}{(6^{16})^{\frac{1}{4}}})\]

OpenStudy (campbell_st):

then simplify

OpenStudy (anonymous):

yeah i said its x =2 was just wanting to see if i was right lol

OpenStudy (campbell_st):

I would have said x = 4 \[\log \frac{6^8}{6^4} = \log(6^4)...then... \log(6^4) = \log(6^x)\]

OpenStudy (anonymous):

Ohhh ops forgot you subtract the power by the number being divided i just divided those 2, thanks for your help goodnight!!